OpenStudy (anonymous):

Solve for y. -1 + y 8 < 15

OpenStudy (anonymous):

can yo write that more neat please

OpenStudy (zepp):

\[-1+8y < 15\] Like this?

OpenStudy (anonymous):

Exactly like that besides the Y is above the 8

OpenStudy (anonymous):

Anything anyone?

OpenStudy (anonymous):

it has to be 1 right guys?

OpenStudy (anonymous):

if it was 2 it would be even and anything highe than 2 wouldnt be less than 15

OpenStudy (zepp):

\[-1 + \frac{y}{8} < 15\] Move -1 to the right \[\frac{y}{8} < 15+1\]

OpenStudy (zepp):

\[\frac{y}{8} < 16\]multiply by \[\frac{y}{8} * 8 < 16*8\]